# A Dirichlet stochastic block model for composition-weighted networks

### v1

@@ Abstract
Network data are observed in various applications where the individual entities of the system interact with or are connected to each other, and very often these interactions are defined by their associated strength or importance. Clustering is a common task in network analysis that involves finding groups of nodes which display similarities in the way they interact with the rest of the network. However, most clustering methods use the strengths of connections between entities in their original form, ignoring the possible differences in the capacities of individual nodes to send or receive edges. This often leads to clustering solutions that are heavily influenced by the nodes’ capacities. One way to overcome this is to analyse the strengths of connections in relative rather than absolute terms, expressing each edge weight as a proportion of the sending (or receiving) capacity of the respective node. This, however, induces additional modelling constraints that most existing clustering methods are not designed to handle. In this work we propose a stochastic block model for composition-weighted networks based on direct modelling of compositional weight vectors using a Dirichlet mixture, with the parameters determined by the cluster labels of the sender and the receiver nodes. Inference is implemented via an extension of the classification expectation-maximisation algorithm that uses a working independence assumption, expressing the complete data likelihood of each node of the network as a function of fixed cluster labels of the remaining nodes. A model selection criterion is derived to aid the choice of the number of clusters. An alternative approach to clustering in composition-weighted networks based on a mapping to the Euclidean space is also provided. The model is validated using a number of simulation studies, assessing the effect of various initialisation strategies on the model’s performance, latent structure recovery, parameter estimation quality and model selection. Furthermore, it is showcased on network data from the Erasmus exchange program and a bike sharing network for the city of London.
Keywords Compositional data, hybrid likelihood, statistical network analysis, stochastic block model, weighted networks
Acknowledgements This work was supported by the Science Foundation Ireland funded Insight Centre for Data Analytics (Grant number 12/RC/2289_P2).
@@ 1 Introduction
All of the aforementioned approaches propose models for the analysis of the weights in their original form. However, in some applications, in particular those modelling flows in the network, such as transportation or trade, the observed edge weights are influenced by the nodes’ capacities to send and receive edges, which can differ significantly across the network. This often results in clustering solutions that mimic the patterns related to the capacities of the sending and receiving nodes, rather than the weights connecting them. Taking as an example the Erasmus programme network presented in Section 6.1, suppose we are interested in grouping European countries based on similarities in students’ preferences for exchange destinations. With the participating countries having very different populations, we expect larger countries, such as Germany, France and the UK, to have larger student populations and so larger volumes of students participating in the programme. These countries are also able to accommodate more incoming students. By contrast, the Nordic countries, like Denmark and Finland, are much smaller in population, and hence they send and receive significantly fewer students. If we were to perform clustering using the raw counts of students going on Erasmus exchange between these countries without taking the differences in population into account, the countries would be roughly divided into groups based on their population sizes, i.e. Germany, France and the UK as one cluster, and Denmark and Finland as another, due to differences in the scale of the weights. Even if the preference profile of the UK students is closer to those of the smaller Nordic countries, the volume of students the country sends is far too large for it to be assigned to the same cluster as Finland and Denmark. Similar issue can arise in flight networks, with small airports only being able to service a limited number of flights per day as opposed to international hubs with multiple runways, or international trade networks, where the value of goods traded is linked to country’s GDP (Melnykov et al.,, 2020). The differences in the magnitudes of edge weights can significantly affect the results, making it more challenging to draw conclusions.
@@
The main challenge associated with the relative nature of edge weights in the network is their interdependence. Proportional data impose a constant-sum constraint, making the standard tools used for networks with independent weights invalid. To appropriately address this constraint, tools from compositional data analysis can be employed. Compositional data describe parts of some whole, such as percentages, probabilities, and proportions, and they can be naturally observed or constructed from other types of data, such as continuous or count data, by dividing individual components by their sum. There are two main families of approaches to compositional data analysis. The first involves mapping the compositions onto the real line, allowing one to use the standard machinery on the transformed data. Log-ratio transformations are a popular choice in the literature due to their sound theoretical properties, and other alternatives, such as standardisation, are also sometimes used in practice (Aitchison,, 1986; Baxter,, 1995). The second approach is based on direct modelling of compositional observations, which often uses the Dirichlet distribution to model the compositions in their original form (Pal and Heumann,, 2022; Tsagris and Stewart,, 2018). Both approaches have their relative benefits, such as the scale-invariance of the log-ratio transformations, or desirable distributional properties that come with direct modelling, and so the choice of the most suitable approach is context dependent. For the discussion of the main methods and challenges in compositional data analysis, we suggest the review paper by Alenazi, (2023).
@@
To the best of our knowledge, there have been no attempts to develop a general model-based clustering framework for composition-weighted networks. In this work, we propose an extension of the stochastic block model that utilises the relative rather than absolute strength of connections, by leveraging ideas from compositional data analysis. We model the set of edge weights from each node using a Dirichlet distribution, the parameters of which are determined by the cluster labels of the sender and each of the receivers. For inference, we consider the hybrid maximum likelihood approach developed by Marino and Pandolfi, (2022), based on Daudin et al., (2008), using a variant of the classification expectation-maximisation (EM) algorithm (Celeux and Govaert,, 1992). To address model selection, an integrated completed likelihood (ICL) criterion is derived (Daudin et al.,, 2008). We also briefly introduce an alternative approach to cluster nodes in composition-weighted networks by making use of the log-ratio transformation that maps compositional data onto the real line (Aitchison,, 1982). The code implementing the modelling framework in R (R Core Team,, 2019) is available on GitHub.
@@ 2 Dirichlet stochastic block model
This section introduces the Dirichlet stochastic block model (DirSBM) for networks with compositional edge weights.
Let $\mathbf{Y}$ be the weighted adjacency matrix of a directed network on $n$ nodes with no self-loops with strictly positive entries, i.e. $y_{ij}>0$ for all $i\neq j$ and $y_{ii}=0$ for all $i$. Define the compositional counterpart of $\mathbf{Y}$, denoted $\mathbf{X}$, whose row entries $\mathbf{x}_{i}$ are given by
@@
i.e. $x_{ij}>0$ for $i\neq j$ and $\sum_{l}x_{il}=1\leavevmode\nobreak\ \forall i$. Note that $x_{ii}=0$ is due to the absence of self-loops in the original weighted network. Denote with $\mathbf{x}_{i}^{*}$ a subvector of $\mathbf{x}_{i}$ that excludes the zero entry $x_{ii}$, and with $\mathbf{X}^{*}$ the matrix formed by row vectors $\mathbf{x}_{i}^{*},i=1,\ldots,n$. In practice, in some applications, the weighted adjacency matrix $\mathbf{Y}$ may be unavailable, with the compositional matrix $\mathbf{X}$ being the only piece of data provided. The Dirichlet stochastic block model with $K$ blocks assumes the following generative process for a network with compositional edge weights $\mathbf{X}^{*}$:
@@
The intuition behind this formulation is as follows: for any sender $i$, the distribution of weights of the edges connecting it to other vertices depends on the cluster memberships of $i$ itself and of all of the receiver nodes $j\neq i$. If the sender $i$ and the receiver $j$ are members of the same cluster $C_{k}$, the corresponding parameter of the Dirichlet distribution is $\alpha_{kk}$, and this parameter value is shared for all such receivers $j\in C_{k}$. If instead the receiver node is in a different cluster, $j\in C_{h}$, the corresponding parameter is $\alpha_{kh}$.
@@ 2.1 Properties of DirSBM
The model statement in Section 2 implies that the distribution of weight vectors conditional on cluster labels in the whole network is shared among the nodes assigned to the same cluster, up to a permutation of the ordering of the nodes, which is assumed to be invariant for simplicity.
To illustrate this, consider, for example, the distribution of compositional weights of node $i\in C_{1}$ and rearrange the order of the dimensions so that they are sorted by cluster assignment index in ascending order for ease of notation. Every vector of compositional weights of node $i$ belonging to cluster $C_{1}$, denoted $\mathbf{x}^{*}_{i\in C_{1}}$, conditional on the cluster assignments of all the remaining $(n-1)$ nodes, is distributed as
@@ 3 Inference
As usual in model-based clustering, inference proceeds by maximizing the marginal log-likelihood of the data $\mathbf{X}$ with respect to the model parameters. However, this quantity involves summing over the set of all possible cluster allocations and is computationally intractable (Daudin et al.,, 2008). A popular algorithm used in model-based clustering is the expectation-maximisation (EM) algorithm (Dempster et al.,, 1977). It is employed to find the maximum of the marginal log-likelihood of the data by introducing latent variables corresponding to cluster labels and defining the joint log-likelihood of the data and the cluster labels, the complete data log-likelihood. Then, inference for the model consists of estimation of parameters and of the latent cluster assignments. The EM algorithm works by iteratively finding the expectation of the complete data log-likelihood with respect to the posterior distribution of the latent variables (E-step) and maximising this expectation with respect to the model parameters (M-step).
In the case of the DirSBM, the complete data log-likelihood is:
@@
Further details are provided in Appendix A.
@@ 3.1 Hybrid log-likelihood
where $\widetilde{\mathbf{Z}}$ is a binary matrix with entries $\tilde{z}_{ik}=\mathbbm{1}\{\tilde{c}_{i}=k\}$ and $\tilde{\alpha}_{j}=\sum_{h=1}^{K}\tilde{z}_{jh}\alpha_{kh}$.
With the complete data hybrid log-likelihood, a working independence assumption allows one to implement a variant of the classification EM algorithm (Celeux and Govaert,, 1992), where the hard partition of the nodes that are not $i$ is used to compute the cluster assignment probability for $i$. The steps of the algorithm are outlined below.
@@
then $\mathbf{c}^{(t)}$ and $\widetilde{\mathbf{Z}}^{(t)}$ are updated. The update of cluster labels happens sequentially for the nodes in the network rather than simultaneously, so any changes in the cluster assignments of nodes before $i$ affect $\tilde{\mathbf{c}}_{-i}$ and hence the cluster label for $i$.
@@
Closed form solution is available for the mixing proportions, so these are updated using
@@
The updates for Dirichlet parameters are not available in closed form, so the set of estimates $\{\hat{\alpha}_{kh}^{(t)}\}_{k,h=1}^{K}$ are found numerically using the R function $optim$ (R Core Team,, 2019). Since the parameters of the Dirichlet distribution are strictly positive, we use the L-BFGS-B optimisation routine, which allows for optimization with parameters having a bounded range (Byrd et al.,, 1995).
@@
$\left|\frac{l^{hyb}(\mathbf{\hat{A}}^{(t)},\boldsymbol{\hat{\theta}}^{(t)})-l^{hyb}(\mathbf{\hat{A}}^{(t-1)},\boldsymbol{\hat{\theta}}^{(t-1)})}{l^{hyb}(\mathbf{\hat{A}}^{(t)},\boldsymbol{\hat{\theta}}^{(t)})}\right|<\epsilon.$ (16)
In this work, $\epsilon=10^{-5}$ was used. More details on the algorithm can be found in Appendix B.

### v2

@@ Abstract
Network data are prevalent in applications where individual entities interact with each other, and often these interactions have associated weights representing the strength of association. Clustering such weighted network data is a common task, which involves identifying groups of nodes that display similarities in the way they interact. However, traditional clustering methods typically use edge weights in their raw form, overlooking that the observed weights are influenced by the nodes’ capacities to distribute weights along the edges. This can lead to clustering results that primarily reflect nodes’ total weight capacities rather than the specific interactions between them. One way to address this issue is to analyse the strengths of connections in relative rather than absolute terms, by transforming the relational weights into a compositional format. This approach expresses each edge weight as a proportion of the sending or receiving weight capacity of the respective node. To cluster these data, a Dirichlet stochastic block model tailored for composition-weighted networks is proposed. The model relies on direct modelling of compositional weight vectors using a Dirichlet mixture, where parameters are determined by the cluster labels of sender and receiver nodes. Inference is implemented via an extension of the classification expectation-maximisation algorithm, expressing the complete data likelihood of each node as a function of fixed cluster labels of the remaining nodes. A model selection criterion is derived to determine the optimal number of clusters. The proposed approach is validated through simulation studies, and its practical utility is illustrated on two real-world networks.
Keywords Compositional data; hybrid likelihood; statistical network analysis; stochastic block model; weighted networks
@@ 1 Introduction
All of the aforementioned approaches propose models for the analysis of the weights in their original form. However, in some applications, in particular those modelling flows in the network (such as transportation or trade), the observed edge weights are influenced by the nodes’ weight capacities to send and receive weights along the edges, which can differ significantly across the network. This often results in clustering solutions that mimic the patterns related to the total weight capacities of the sending and receiving nodes, rather than the individual weights connecting them. Note that by the term weight capacity we mean the total volume of the weights each node can send or receive along the edges, which differs from both the total number of edges a node can send or receive and its degree (see for example Karrer and Newman,, 2011). Taking as an example the Erasmus programme network presented in Section 6.1, suppose we are interested in grouping European countries based on similarities in students’ preferences for exchange destinations. With the participating countries having very different populations, we expect larger countries, such as Germany, France and the UK, to have larger student populations and hence larger volumes of students participating in the programme. These countries are also able to accommodate more incoming students. By contrast, the Nordic countries, like Denmark and Finland, are much smaller in population, and hence they send and receive significantly fewer students. If we were to perform clustering using the raw counts of students going on Erasmus exchange between these countries without taking the differences in population into account, the countries would be roughly divided into groups based on their population sizes, i.e. Germany, France and the UK as one cluster, and Denmark and Finland as another, due to differences in the scale of the weights. Even if the preference profile of the UK students is closer to those of the smaller Nordic countries, the volume of students the country sends is far too large for it to be assigned to the same cluster as Finland and Denmark. Similar issue can arise in flight networks, with small airports only being able to service a limited number of flights per day as opposed to international hubs with multiple runways, or international trade networks, where the value of goods traded is linked to country’s GDP (Melnykov et al.,, 2020). The differences in the magnitudes of edge weights can significantly affect the results, making it more challenging to draw conclusions.
@@
The main challenge associated with the relative nature of edge weights in the network is their interdependence. Proportional data impose a constant-sum constraint, making the standard tools used for networks with independent weights invalid. To appropriately address this constraint, tools from compositional data analysis can be employed. Compositional data describe parts of some whole, such as percentages, probabilities, and proportions, and they can be naturally observed or constructed from other types of data, such as continuous or count data, by dividing individual components by their sum. There are two main families of approaches to compositional data analysis. The first involves mapping the compositions onto the real line, allowing one to use the standard machinery on the transformed data. Log-ratio transformations are a popular choice in the literature due to their sound theoretical properties, and other alternatives, such as standardisation, are also sometimes used in practice (Aitchison,, 1986; Baxter,, 1995). The second approach is based on direct modelling of compositional observations, which often uses the Dirichlet distribution to model the compositions in their original form (Pal and Heumann,, 2022; Tsagris and Stewart,, 2018). Both approaches have their relative benefits, such as the scale-invariance of the log-ratio transformations, or desirable distributional properties that come with direct modelling, therefore, the choice of the most suitable approach is context dependent. For a discussion of the main methods and challenges in compositional data analysis, we suggest the review paper by Alenazi, (2023).
@@
To the best of our knowledge, there have been no attempts to develop a general model-based clustering framework for composition-weighted networks. In this work, we propose an extension of the stochastic block model for dense networks of moderate size (in terms of the number of nodes) that utilises the relative rather than absolute strength of connections, by leveraging ideas from compositional data analysis. We model the set of edge weights from each node using a Dirichlet distribution, the parameters of which are determined by the cluster labels of the sender and each of the receivers. For inference, we consider the hybrid maximum likelihood approach developed by Marino and Pandolfi, (2022), using a variant of the classification expectation-maximisation (EM) algorithm (Celeux and Govaert,, 1992). To address model selection, an integrated completed likelihood (ICL) criterion is derived (Daudin et al.,, 2008). We also briefly introduce an alternative approach to cluster nodes in composition-weighted networks by making use of the log-ratio transformation that maps compositional data onto the real line (Aitchison,, 1982). The code implementing the modelling framework in R (R Core Team,, 2024) is available on GitHub.
@@ 2 Dirichlet stochastic block model
This section introduces the Dirichlet stochastic block model (DirSBM) for networks with compositional edge weights. Let $\mathbf{Y}$ be the weighted adjacency matrix of a directed network on $n$ nodes with no self-loops with strictly positive entries, i.e. $y_{ij}>0$ for all $i\neq j$ and $y_{ii}=0$ for all $i$. Define the compositional counterpart of $\mathbf{Y}$, denoted $\mathbf{X}$, whose row entries $\mathbf{x}_{i}$ are given by
@@
i.e. $x_{ij}>0$ for $i\neq j$ and $\sum_{l}x_{il}=1\leavevmode\nobreak\ \forall i$. Note that $x_{ii}=0$ due to the absence of self-loops in the original weighted network. Denote with $\mathbf{x}_{i}^{*}$ a subvector of $\mathbf{x}_{i}$ that excludes the zero entry $x_{ii}$, and with $\mathbf{X}^{*}$ the matrix formed by row vectors $\mathbf{x}_{i}^{*},i=1,\ldots,n$. In practice, in some applications, the weighted adjacency matrix $\mathbf{Y}$ may be unavailable, with the compositional matrix $\mathbf{X}$ being the only piece of data provided. The Dirichlet stochastic block model with $K$ blocks assumes the following generative process for a network with compositional edge weights $\mathbf{X}^{*}$:
@@
The intuition behind this formulation is as follows: for any sender $i$, the distribution of weights of the edges connecting it to other vertices depends on the cluster memberships of $i$ itself and of all of the receiver nodes $j\neq i$. If the sender $i$ and the receiver $j$ are members of the same cluster $C_{k}$, the corresponding parameter of the Dirichlet distribution is $\alpha_{kk}$, and this parameter value is shared for all such receivers $j\in C_{k}$. If instead the receiver node is in a different cluster, $j\in C_{h}$, the corresponding parameter is $\alpha_{kh}$. Since the model does not include any node-specific parameters, node-level heterogeneity is not accounted for in this framework.
The proposed model aims to describe the clustering structure of directed fully-connected networks, with the edge weights being expressed as proportions sent by the nodes (i.e. the sum of outgoing weights for each node is one). Note that the decision to use sending proportions rather than receiving proportions (i.e. where the sum of incoming weights for each node is one) is arbitrary. In fact, the modelling of receiving proportions can be achieved by working with the transpose of $\mathbf{Y}$ instead of $\mathbf{Y}$ itself.
Although DirSBM represents a novel approach for model-based clustering of proportion-weighted networks, it is not the first model that aims to account for the node capacities in the network. For example, in Karrer and Newman, (2011), the authors propose the degree-corrected stochastic block model (DC-SBM), which allows better description of real-world networks with degree heterogeneity by introducing node-specific parameters. DirSBM and DC-SBM, however, serve different modeling purposes and consider node capacities differently. DC-SBM accounts for variations in the nodes’ capacities to connect to other nodes in the network, allowing nodes assigned to the same cluster to have distinct degree distributions. DirSBM, on the other hand, does not aim to address the node degree heterogeneity. Instead, it focuses on discounting the effects of nodes’ weight capacities, that is, the total weights each node can distribute along its edges, which can vary significantly across the network. DirSBM achieves this by representing relational data in a compositional format, thereby neutralising the impact of the total weight a node sends to (or receives from) others. This approach allows clustering to be informed by relative proportions. Moreover, DirSBM is also designed for fully-connected or highly dense networks whose degree distributions do not typically follow a power-law.
In addition to modeling proportional rather than absolute weights, DirSBM is also distinguished from existing SBMs (particularly weighted SBMs) by its relaxation of the independence assumption between edges in the network, given the cluster labels of their end nodes. For instance, Aicher et al., (2014) introduces a general weighted stochastic block model that balances the importance of edge presence and edge weights, with the latter drawn from an exponential family of distributions. However, this model assumes conditional independence: both binary edges and edge weights are independent, given the cluster labels, and edge weights are independent of the presence of binary edges. Similarly, Ng and Murphy, (2021) proposes a hierarchical model in which gamma-distributed weights are attributed only to existing binary edges. However, like Aicher et al., (2014), it assumes that binary edges and edge weights are conditionally independent, conditional on the cluster labels of their end nodes. To the best of our knowledge, DirSBM is a novel approach in its relaxation of the independence assumption between the edges.
@@ 2.1 Properties of DirSBM
The model statement in Section 2 implies that the distribution of weight vectors conditional on cluster labels in the whole network is shared among the nodes assigned to the same cluster, up to a permutation of the ordering of the nodes, which is assumed to be invariant for simplicity. To illustrate this, consider, for example, the distribution of compositional weights of node $i\in C_{1}$ and rearrange the order of the dimensions so that they are sorted by cluster assignment index in ascending order for ease of notation. Every vector of compositional weights of node $i$ belonging to cluster $C_{1}$, denoted $\mathbf{x}^{*}_{i\in C_{1}}$, conditional on the cluster assignments of all the remaining $(n-1)$ nodes, is distributed as
@@ 2.2 Interpretation of parameters
2.3 Networks with zero-weighted edges
As stated earlier, DirSBM is developed for fully connected networks characterized by weighted adjacency matrices with strictly positive entries. However, as shown in the applications presented in Section 6, even the most dense real-world networks exhibit some percentage of zeros in the weighted adjacency matrix. These zeros in the adjacency matrix may represent either the absence of edges between certain pairs of nodes, or the presence of edges with zero weight. The distinction between the two often affects the modelling process, as the former relates to network sparsity and the latter does not. An edge is absent in the network where a connection between the nodes is not possible or has not been observed. A zero-weighted edge, on the other hand, is defined as a present connection between the nodes that is attributed with a zero weight or strength. For instance, in the context of the Erasmus network from Section 6.1, exchange between any pair of countries is possible, therefore zeros in the raw data matrix can be viewed as zero weights.
Since DirSBM is designed for clustering fully-connected networks, modelling network sparsity is outside the scope of this work. Consequently, we treat any zeros in the weighted adjacency matrix as zero-weighted edges rather than absent links. To address zero-weighted edges, we adopt a standard procedure from compositional data analysis, where zeros present unique challenges and remain an active area of research, as standard tools are not designed to handle zero-valued compositional parts (Filzmoser et al.,, 2018). A common approach is to replace zeros with a small constant value, as this method is simple and generally effective. Therefore, to address zero-weighted edges in the data, a small value $\epsilon$ is added to the raw data before converting it to compositions. The final results are typically robust relative to the magnitude of the added constant since the small value is added to the raw data rather than to the final compositions. In Section 5.5 we assess the impact of the proportion of zero-weighted edges in the network on parameter estimation and clustering performance of DirSBM.
@@ 3 Inference
As usual in model-based clustering, inference proceeds by maximizing the marginal log-likelihood of the data $\mathbf{X}$ with respect to the model parameters. However, this quantity involves summing over the set of all possible cluster allocations and is computationally intractable (Daudin et al.,, 2008). A popular algorithm used in model-based clustering is the expectation-maximisation (EM) algorithm (Dempster et al.,, 1977). It is employed to find the maximum of the marginal log-likelihood of the data by introducing latent variables corresponding to cluster labels and defining the joint log-likelihood of the data and the cluster labels, the complete data log-likelihood. Then, inference for the model consists of estimation of parameters and of the latent cluster assignments. The EM algorithm works by iteratively finding the expectation of the complete data log-likelihood with respect to the posterior distribution of the latent variables (E-step) and maximising this expectation with respect to the model parameters (M-step). In the case of the DirSBM, the complete data log-likelihood is:
@@
Further details are provided in A.
@@ 3.1 Hybrid log-likelihood
where $\widetilde{\mathbf{Z}}$ is a binary matrix with entries $\tilde{z}_{ik}=\mathbbm{1}\{\tilde{c}_{i}=k\}$ and $\tilde{\alpha}_{j}=\sum_{h=1}^{K}\tilde{z}_{jh}\alpha_{kh}$. With the complete data hybrid log-likelihood, a working independence assumption allows one to implement a variant of the classification EM algorithm (Celeux and Govaert,, 1992), where the hard partition of the nodes that are not $i$ is used to compute the cluster assignment probability for $i$. The steps of the algorithm are outlined below.
@@
then $\mathbf{c}^{(t)}$ and $\widetilde{\mathbf{Z}}^{(t)}$ are updated. The update of cluster labels happens sequentially for the nodes in the network rather than simultaneously, therefore any changes in the cluster assignments of nodes before $i$ affect $\tilde{\mathbf{c}}_{-i}$ and hence the cluster label for $i$.
@@
Closed form solution is available for the mixing proportions, which are updated using
@@
The updates for Dirichlet parameters are not available in closed form, and the set of estimates $\{\hat{\alpha}_{kh}^{(t)}\}_{k,h=1}^{K}$ are found numerically using the R function $optim$ (R Core Team,, 2024). Since the parameters of the Dirichlet distribution are strictly positive, we use the L-BFGS-B optimisation routine, which allows for optimization with parameters having a bounded range (Byrd et al.,, 1995).
@@
$\left|\frac{l^{hyb}(\mathbf{\hat{A}}^{(t)},\boldsymbol{\hat{\theta}}^{(t)})-l^{hyb}(\mathbf{\hat{A}}^{(t-1)},\boldsymbol{\hat{\theta}}^{(t-1)})}{l^{hyb}(\mathbf{\hat{A}}^{(t)},\boldsymbol{\hat{\theta}}^{(t)})}\right|<\varepsilon.$ (16)
In this work, $\varepsilon=10^{-5}$ was used. More details on the algorithm can be found in B.